OpenStudy (anonymous):

Evaluate 7x - (2x)^2 + x^3 for x = -2.

OpenStudy (anonymous):

Just plug in -2 for every x

OpenStudy (anonymous):

7(-2)-(2(-2))^2+(-2)^3

OpenStudy (anonymous):

7(-2) -(-4)^2 + (-2)^3 = -14 - 16 -8 = -38

OpenStudy (anonymous):

plug -2 everywhere there is an x

OpenStudy (anonymous):

7(-2) - [2(-2)]^2 + (-2)^3 -14 - [-4]^2 -8 -14 - 16 - 8 -38